OpenStudy (anonymous):

write an equation in standard form for the line (-3,14) and (-6,9)

OpenStudy (ash2326):

@charitynelson If we have two point (x1, y1) and (x2, y2) equation of line passing through any point (x, y) is given as \[\frac{y-y_2}{x-x_2}=\frac{y-y_1}{x-x_1}\] Can you plugin (x1, y1)= (-3, 14) and (x2, y2)= (-6, 9) here?

OpenStudy (anonymous):

woah im lost

OpenStudy (anonymous):

is that how its set up jw

OpenStudy (ash2326):

just plugin the values, substitute for x1, y1, x2, and y2

OpenStudy (anonymous):

so would it be -3--6 and 14-9

OpenStudy (ash2326):

ok, I'll show \[\frac{y-y_2}{x-x_2}=\frac{y_2-y_1}{x_2-x_1}\] (x1, y1)= (-3, 14) and (x2, y2)= (-6, 9) \[\frac{y-(9)}{x-(-6)}=\frac{9-(14)}{(-6)-(-3)}\] Can you simplify this ?

OpenStudy (anonymous):

for the first one i got 3/2 and the second one i got -5/-3

OpenStudy (ash2326):

@charitynelson notice that on the left side we have variables x and y, we can't reduce it further. However we can reduce the right side so we have \[\frac{y-9}{x+6}=\frac{-5}{-3}\] Can you cross multiply and simplify ?

OpenStudy (anonymous):

are we cross multiplying the fraction and do i flip the -5/-3

OpenStudy (ash2326):

yes. cross multiplying and yes you have to flip the fraction when you take it to the left side

OpenStudy (anonymous):

can u show me what u mean by bringing to the left side because the wya i have it wrote idk just doesnt seem right

OpenStudy (ash2326):

just a min

OpenStudy (anonymous):

thank you

OpenStudy (ash2326):

We can cancel both the negatives on left \[\frac{y-9}{x+6}=\frac{5}{3}\] Let's cross multiply \[3(y-9)=5(x+6)\] \[3y-27=5x+30\] Can you simplify this from here? bring variables to one side and constants on other

OpenStudy (anonymous):

thank you sooo much and i had another question its quick he wants me to find slop of a line perpendicular to the graph of x-y=4

OpenStudy (ash2326):

@charitynelson can you find slope of this line x-y=4?

OpenStudy (anonymous):

no my teacher never explained well enough i know you use y = mx+b

OpenStudy (ash2326):

yes, just bring the equation to this form \[y=mx+b\]

OpenStudy (anonymous):

so how do we plug that equation in

OpenStudy (ash2326):

it's easy, change it to that form \[x-y=4\] \[x-4=y\] or \[y=x-4\] now you can find m which is slope

OpenStudy (anonymous):

this is the part the trows me off my teacher gives me the answer which is one but never explained

OpenStudy (ash2326):

just compare both \[y=mx+b\] \[y=1x-4\] m=1 b=-4

OpenStudy (anonymous):

so for perpendicular how did he get -1

OpenStudy (ash2326):

If two lines have slopes m1 and m2 and they are perpendicular, then \[m_1\times m_2=-1\] here m1 is 1 we need to find m2 \[1\times m_2=-1\] you can find slope of perpendicular line from this

OpenStudy (anonymous):

so how does m1 fit into beling -1 do we just subtract it or

OpenStudy (ash2326):

no m1=1, which is the slope of given line but m2, slope of perpendicular line =-1

OpenStudy (anonymous):

so its the opposit of what ever slope is

OpenStudy (ash2326):

if you have 2, then perpendicular line's slope =-1/2

OpenStudy (anonymous):

so say for instnace it says to find slope of 2x-y=4

OpenStudy (anonymous):

then would the slope be 1/2

OpenStudy (anonymous):

or 2

OpenStudy (ash2326):

Slope of this line, you have to compare from y=mx+c \[y=mx+c\] \[y=2x-4\] m=2 c=-4

OpenStudy (anonymous):

this is so confusing

OpenStudy (ash2326):

it's just comparison, forget about reversing I have \[y=2x+4\] I'll compare it yo y=mx+c m=2 and c=4 so slope of this line =2

OpenStudy (anonymous):

ok and then now the perpendicular part

OpenStudy (ash2326):

for that \[2\times m_2=-1\] m2=-1/2 slope of line perpendicular to original line

OpenStudy (anonymous):

oh i see it now ahhaha wow

OpenStudy (ash2326):

cool

OpenStudy (anonymous):

do u have time to help me with a more complicated problem

OpenStudy (ash2326):

@charitynelson I have to go :( Sorry. I'll try to come later and help.